OpenStudy (anonymous):

What is the circumference of a circle with a diameter of 6?

OpenStudy (anonymous):

divide 6 by 2

OpenStudy (anonymous):

That gives you R Then go @r*pi

OpenStudy (anonymous):

so 3*2*pi

OpenStudy (anonymous):

Do they want the exact answer>

OpenStudy (anonymous):

Or an estimate?

OpenStudy (anonymous):

Estimate: 18.84 Exact: 6pi

OpenStudy (anonymous):

^interesting ;)

OpenStudy (anonymous):

The equation for circumference is \[\pi d\] where d = diameter so you have \[6 \pi\] or roughly 18

OpenStudy (wwe123):

2*pi*r 2*22/7*6/2 132/7 18.85

OpenStudy (anonymous):

Like I said.

OpenStudy (jiteshmeghwal9):

\[{\color{red}{Circumfrence \space of \space a \space circle=\pi \times diameter}}\]

OpenStudy (jiteshmeghwal9):

\[\color{green}{{3.14}\times6}\]

OpenStudy (anonymous):

\[C = 2\pi r\]

OpenStudy (anonymous):

WE OBVIOUSLY ALREADY HAVE THE ANSWER SO STOP POSTING>

OpenStudy (anonymous):

The answer will depend on the value of pi given you

OpenStudy (anonymous):

Lol. Sorry for the rage.

OpenStudy (anonymous):

you don't expect to get the same answer when pi = 22/7 as when pi = 3.14 or when pi = 3.142

OpenStudy (anonymous):

OF course not. But generally in highschool courses they have you use just 3.14 In college they have you use: 3.14159 or 3.1416

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

d = 2r \[\rightarrow C = \pi d\]